OpenStudy (anonymous):

find the difference quotient of f; that is, find f(x+h)-f(x)/h, h cannot equal 0, a. f(x)=3x^2-2x+6

OpenStudy (anonymous):

apply l hospital rule

OpenStudy (anonymous):

I don't knwo what that means?

OpenStudy (anonymous):

ok no problem just express the function u want like fx-fx+h/h

OpenStudy (anonymous):

this is what i have so far f(x+h)=3(x+h)^2-2(x+h)+6 I don't know where to go from there

OpenStudy (anonymous):

take x common out of it and expand it with binomial just taking ist two terme out of it

OpenStudy (anonymous):

so just 3^2-2+6? since i took out the common (x+h)

OpenStudy (anonymous):

i m working it out plz wait

OpenStudy (anonymous):

okay. thank you

OpenStudy (anonymous):

u hav taken x+h common take x common

OpenStudy (anonymous):

by the way i hav a short cut just differentiate it

OpenStudy (anonymous):

I realized i wrote the problem wrong. its find the expression for f(x+h) by substituting x+h for each x in the expression for f(x) a. f(x+h)=3x^2-2x+6

OpenStudy (anonymous):

u must hav got answer as 6x-2

OpenStudy (anonymous):

i don't have anything yet, i'm unsure of how to solve it

OpenStudy (anonymous):

see follow the steps just arrange the whole function in the form fx+h-fx/h

OpenStudy (anonymous):

then take x common from the term which hav x+h and then expand it with binomial to get rid from the degrees